OpenStudy (owlcoffee):

A 3rd degree function "f" satisfies the following: f'(2)=-1 f''(4)=0 Calculate f'(6).

imqwerty (imqwerty):

f' is the derivative of f right?

OpenStudy (owlcoffee):

yeah, I believe so. I found this and I am skeptic if the given information is sufficient to deduce that...

OpenStudy (anonymous):

The key is the phrase "third degree function", which I assume means "third degree polynomial". If that's the case, then the information is sufficient.

OpenStudy (owlcoffee):

Well, the derivative of a 3rd degree polynomial becomes a 2nd degree polynomial, which would imply that two, one or none extrema should exist.

OpenStudy (anonymous):

Because there is never any reference to f itself, we may as well consider a second degree function g = f' such that g(2) = -1 g'(4) = 0 That implies that g has either a minimum or a maximum at x=4. In either case, g(6) should be equal to g(2), so the answer to the question is g(6) = f'(6) = -1.

imqwerty (imqwerty):

ok its a 3rd degree function so lets say\[f(x)=ax^3+bx^2+cx+d\] now\[f'(x)=3ax^2+2bx+c\]\[f'(2)=3a(2^2)+2b(2)+x=-1 \] so u have 1st equation as-> \[12a+4b+c=-1\] now\[f''(x)=6ax+2b\]\[f''(4)=6a(4)+2b=0\]\(24a+2b=0\) -->2nd equation now we need to find f''(6) \[f'(6)=108a+12b+c\] now just try to add eq 1 with 4 times of equation 2 :D

OpenStudy (owlcoffee):

yeah, I found that as a possible process of solution, I'm guessing that finding the function is the way to go, I thought that there could be any generalizations possible from this (I haven't touched derivatives in a while now).

imqwerty (imqwerty):

\(\Huge\ddot\smile\)

OpenStudy (owlcoffee):

By the way, I think it becomes a problem, since it'll be two equations with three variables.

imqwerty (imqwerty):

We need to find ---->\(f'(6)=108a+12b+c\) 1st equation-------->\(-1=12a+4b+c\) 2nd equation------->\(0=24a+2b\) adding the eq 1 to 4 times of equation 2 u get this-> \(-1+4(0)=12a+4b+c+4(24a+2b)\) \(-1=108a+12b+c\) but \(f'(6)=108a+12b+c\) so\(f'(6)=108a+12b+c=-1\)

OpenStudy (owlcoffee):

Ah, I see, you worked the algebra to make their expressions fit, and thus one being equal to the other.

imqwerty (imqwerty):

e.e yes!

Parth (parthkohli):

If we let\[f''(x) = k(x - 4)\]and step backwards, then we get \(f'(x) = kx^2/2 - 4kx + c\). Now \(f'(2) =-1\) so \(c = 6k-1\). In this case, we get \(f'(x) = kx^2/2 - 4kx + 6k-1\) and so \(f'(6) = -1\).